# HopRAG: Multi-Hop Reasoning for Logic-Aware Retrieval-Augmented Generation

### v1

@@ 4.2 Main Results
The main results are presented in Table 1 and 2. We observe that almost in all the settings our method gives the best performance, with exceptions on HotpotQA when compared against SiReRAG and 2WikiMultiHopQA against HippoRAG. Overall, HopRAG achieves approximately 76.78% higher than dense retriever (BGE), 9.94% higher than RAPTOR, 3.08% higher than HippoRAG, 1.11% higher than SiReRAG. This illustrates the strengths of HopRAG in capturing both textual similarity and logical relations for handling multi-hop QA.
Specifically, unstructured RAG baselines sparse retriever (BM25) and dense retriever (BGE) yield unsatisfactory results since they rely solely on similarity. Since GraphRAG only considers relevance among entities instead of similarity for graph search, and RAPTOR mainly focuses on the hierarchical logical relations among passages but cannot capture other kinds of relevance, both of them are more suitable for query-focused summarization but they are not the most competitive method for multi-hop QA tasks, as also reported in (Zhang et al., 2024).
@@
Table 5: We conduct an ablation study on whether to use a reasoning model (GPT-4o-mini) during traversal with GPT-3.5-turbo as the inference model and top 20 passages. We compare 4 scenarios including sparse retriever (BM25), dense retriever (BGE), HopRAG (non-LLM) and HopRAG (GPT-4o-mini) and report both the answer and the retrieval metrics, where the best score is in bold and the second best is underlined.
In terms of HippoRAG, it prioritizes relevance signals such as vertices with the most edges and does not explicitly model similarity in its design while our design HopRAG directly integrates similarity with logical relations when constructing edges. Although HopRAG only outperforms SiReRAG by a small margin in the scenario with top 20 candidate passages, our general graph structure does not introduce additional summary and proposition aggregate nodes and can facilitate efficient multi hops for faster retrieval compared with SiReRAG. In the discussion, we will show that HopRAG can achieve competitive results with smaller context length. Besides quantitative scores, we also demonstrate a case study in Appendix A.6.
@@ 4.3 Ablations and Discussion
To confirm the robustness of HopRAG and provide more insights, we vary hyperparameters $top_{k}$, $n_{hop}$ and conduct ablation studies on whether to use a reasoning model during traversal.
@@
To show our efficiency in faster hopping from indirectly relevant passages to truly relevant ones, we test the robustness by evaluating both the QA and retrieval performance on GPT-3.5-turbo with smaller $top_{k}$, as is shown in Table 3. From the results we find that even with top 12 candidates, the QA performance of HopRAG is still comparable with HippoRAG or RAPTOR with 20 candidates, which emphasizes the advantage of our graph traversal design on efficient hopping to retrieve more information within a limited context length. Meanwhile, we also observe that as $top_{k}$ and context length increase, the retrieval F1 score gradually decreases due to the inclusion of excessive redundant information. Conversely, the answer quality generally improves, attributed to GPT-3.5-turbo’s strong capability in processing and reasoning over extended contexts, with only one exception in the MuSiQue dataset where the answer Exact Match (EM) score is slightly higher with top 8 candidates compared to top 12. This suggests that excessive redundancy in the context can, to some extent, hinder the LLM’s ability to identify truly useful information during QA inference.
@@
To assess the effect of the hyperparameter $n_{hop}$ in reasoning-augmented graph traversal, we varied $n_{hop}$ from 1 to 4 and evaluated the corresponding retrieval performance and cost of graph traversal, which is represented by the total number of calling LLM for each vertex in the queue in each round of hop. The results shown in Table 4 indicate that as $n_{hop}$ increases, retrieval performance tends to improve, as more vertices are covered during traversal for reasoning and pruning. However, the expense of calling LLM also increases with $n_{hop}$, creating a trade-off between performance and cost. To find the ideal $n_{hop}$, we notice that as $n_{hop}$ increases, the number of new vertices that require LLM reasoning in the queue in each hop decays quickly, as different vertices may hop to the same important vertex, making the actual queue length of each hop less than $top_{k}$. Specifically, the average queue length in the fourth hop is only 2.60 while in the fifth hop it is only 1.23 so there is no absolute need for one more hop. Based on these findings, we believe that 4 is the ideal $n_{hop}$ for the three datasets, as is what we set for $n_{hop}$ in Table 1 and 2. We also evaluate the answer performance as $n_{hop}$ varies and show the overall results in the appendix A.7.
Ablation on Reasoning Model
In order to generalize our method to the scenario with no budget of calling LLM during retrieval, we provide a graph traversal version based purely on logical relations and textual similarity. Specifically, we replace the Reasoning phase in Algorithm 1 with similarity matching where we directly choose the most similar edge ${e_{j,k}}$ with the query $q$ and hop to $v_{k}$ to extend $C_{queue}$. The results are shown in Table 5. From the results we observe that even without using the reasoning ability of LLM in the graph traversal, HopRAG can achieve 45.84% higher than BM25 and 25.43% higher than dense retriever (BGE), which shows that our graph-structured index can fully capture the textual similarity and logical relations among different vertices to assist in logic-aware retrieval. But as is discussed in Section 1, high similarity doesn’t necessarily mean logical relevance so the introduction of reasoning ability from LLM is still irreplaceable and it can achieve about 45.78% higher average score than non-LLM version. Non-LLM graph traversal is suitable for scenarios that require efficiency, while reasoning-augmented graph traversal is designed for QA tasks that require high accuracy.
@@ 5 Conclusion
Extensive experiments on multi-hop QA benchmarks, i.e. MuSiQue, 2WikiMultiHopQA, and HotpotQA, demonstrate that HopRAG outperforms conventional RAG systems and state-of-the-art baselines. Specifically, HopRAG achieved over 76.78% higher answer accuracy and 65.07% improved retrieval F1 score compared to dense retrievers. It highlights the effectiveness of integrating logical reasoning into the retrieval module. Moreover, ablation studies provide insights into the sensitivity of hyperparameters such as top-k and hop number, revealing trade-offs between retrieval performance and computational costs.
HopRAG paves the way toward reasoning-driven knowledge retrieval. Future work involves scaling HopRAG to broader domains beyond QA tasks; optimizing the indexing and traversal strategies to handle even more complex scenarios with lower computation costs.
6 Limitations
@@ A.1 Symbols
$P$ passage corpus $Q_{i}^{+}$ set of out-coming quesitions for $p_{i}$
$p$ passage $q_{i,j}^{+}$ the j-th out-coming quesition for $p_{i}$
$q$ query $Q_{i}^{-}$ set of in-coming quesitions for $p_{i}$
$C$ retrieval context $q_{i,j}^{-}$ the j-th in-coming quesition for $p_{i}$
@@
$\mathcal{E}$ set of directed edges $v_{i,j}^{+}$ embedding for $q_{i,j}^{+}$
@@
$k_{q}$ keywords for $q$ $v_{i,j}^{-}$ embedding for $q_{i,j}^{-}$
$v_{q}$ embedding for $q$ $R^{+}_{i}$ set of out-coming triplets
$C^{\prime}$ counter of vertices during traversal $r_{i,j}^{+}$ out-coming triplet for $q_{i,j}^{+}$
@@ A.3 Metrics
$F_{1_{\text{retrieval}}}=\frac{2\cdot P\cdot R}{P+R}$
@@ A.4 Settings
To avoid semantic loss by chunking the documents at a fixed size, we chunk each document in a way corresponding to the supporting facts of each dataset. Specifically, we chunk each document in HotpotQA and 2WikiMultiHopQA by sentence since the smallest unit of these two datasets’ supporting facts is a sentence. To get embedding representation for each chunk, we use bge-base model. To extract keywords, we use the part-of-speech tagging function of Python package PaddleNLP to extract and filter entities. In our method, we use the Neo4j graph database to store vertices and build edges. When building edges we employ prompt engineering technique to instruct the LLM to generate an appropriate number of questions for each vertex to cover its information, with a minimum requirement of at least 2 in-coming questions and 4 out-coming questions. To prevent the graph structure from becoming overly complex and dense, we retain only $O(n\cdot\log(n))$ edges, where $n$ is the number of vertices. We use GPT-4o-mini for reasoning-augmented graph traversal, GPT-4o and GPT-3.5-turbo for inference with 2048 max tokens and 0.1 temperature.
@@ A.5 Prompts
(a) Demonstration of our graph structure.
@@
Figure 7: Prompt for reasoning-augmented graph traversal.
Figure 8: Demonstration of one edge between two vertices.
@@ A.6 Case Study
We demonstrate the graph structure in Figure 6(a), one example edge with two vertices in Figure 8. Using the query "Donnie Smith who plays as a left back for New England Revolution belongs to what league featuring 22 teams?" as an example we conduct a qualitative analysis. For this multi-hop question (correct answer: Major League Soccer), the HotpotQA corpus contains three relevant sentences: (1) "Donald W. \"Donnie\" Smith (born December 7, 1990 in Detroit, Michigan) is an American soccer player who plays as a left back for New England Revolution in Major League Soccer. "; (2) "Major League Soccer (MLS) is a men’s professional soccer league, sanctioned by U.S. Soccer, that represents the sport’s highest level in both the United States and Canada. " and (3) "The league comprises 22 teams in the U.S. and 3 in Canada."
@@
A.7 Discussion results on $n_{hop}$
@@
Table 8: We test the effect of hyperparameter $n_{hop}$ on HopRAG using GPT-3.5-turbo on multiple datasets with top 20 passages. We vary $n_{hop}$ from 1 to 4 and report both the answer and retrieval metrics here. For answer metrics, we report the answer EM and F1 score; For retrieval metrics, we report the F1 score and average number of calling LLM during traversal to measure the cost (the lower, the better). The best score is in bold and the second best is underlined.

### v2

@@ 4.2 Main Results
The main results are presented in Table 1 and 2. We observe that almost in all the settings HopRAG gives the best performance, with exceptions on HotpotQA when compared against SiReRAG and 2WikiMultiHopQA against HippoRAG. Overall, HopRAG achieves approximately 76.78% higher than dense retriever (BGE), 48.62% higher than query decomposition, 36.25% higher than reranking (BGE), 9.94% higher than RAPTOR, 3.08% higher than HippoRAG, 1.11% higher than SiReRAG. This illustrates the strengths of HopRAG in capturing both textual similarity and logical relations for handling multi-hop QA.
Specifically, BM25, BGE and BGE with query decomposition yield unsatisfactory results since they rely solely on similarity, and BGE with reranking cannot capture logical relevance among candidates. Since GraphRAG considers relevance among entities instead of similarity for graph search, and RAPTOR focuses on the hierarchical logical relations among passages but cannot capture other kinds of relevance, both of them are more suitable for query-focused summarization but not the most competitive method for multi-hop QA tasks, as also reported in (Zhang et al., 2024).
@@
Table 5: We conduct an ablation study on the reasoning model during traversal with GPT-3.5-turbo as the inference model and top 20 passages. We compare 5 scenarios including sparse retriever (BM25), dense retriever (BGE), HopRAG (non-LLM), HopRAG (Qwen2.5-1.5B-Instruct) and HopRAG (GPT-4o-mini) and report both the answer and the retrieval metrics, where the best score is in bold and the second best is underlined.
In terms of HippoRAG, it prioritizes relevance signals such as vertices with the most edges and does not explicitly model similarity while our design HopRAG directly integrates similarity with logical relations when constructing edges. Although HopRAG only outperforms SiReRAG by a small margin in the scenario with top 20 candidate passages, our general graph structure does not introduce additional summary and proposition aggregate nodes and can facilitate efficient graph traversal for faster retrieval compared with SiReRAG. In the discussion, we will show that HopRAG can achieve competitive results with a smaller context length. Besides quantitative scores, we also demonstrate a case study in Appendix A.6 comparing HopRAG and GraphRAG.
@@ 4.3 Ablations and Discussion
To confirm the robustness of HopRAG and provide more insights, we vary $top_{k}$, $n_{hop}$ and conduct ablation studies on traversal model.
@@
To show our efficiency in faster hop from indirectly relevant passages to truly relevant ones, we test the robustness by evaluating both the QA and retrieval performance on GPT-3.5-turbo with smaller $top_{k}$, as is shown in Table 3. From the results, we find that even with top 12 candidates, the QA performance of HopRAG is still comparable to that of HippoRAG or RAPTOR with 20 candidates, which highlights the effectiveness of our graph traversal design in efficiently retrieving more information within a limited context length. Meanwhile, we also observe that as $top_{k}$ increases, the retrieval F1 score gradually decreases due to the inclusion of excessive redundant information. Conversely, the answer quality generally improves, attributed to GPT-3.5-turbo’s strong capability in processing and reasoning over extended contexts, with only one exception in the MuSiQue dataset.
@@
To assess the effects of the hyperparameter $n_{hop}$ on reasoning-augmented graph traversal, we vary $n_{hop}$ from 1 to 4 and evaluate the corresponding retrieval performance and cost, which is measured by the total number of LLM calls during graph traversal. The results shown in Table 4 indicate that as $n_{hop}$ increases, retrieval performance tends to improve, as more vertices are visited during traversal for reasoning and pruning. However, the expense and latency from calling LLM also increase with $n_{hop}$, creating a trade-off between performance and cost. We notice that as $n_{hop}$ increases, the number of new vertices in $C_{queue}$ requiring LLM reasoning decays rapidly. Since different vertices may hop to the same important vertex, the actual queue length in each round of hop is less than $top_{k}$. Specifically, the average queue length is 2.60 in the fourth round and 1.23 in the fifth round, suggesting that for the three datasets, the local area in the graph structure can be largely explored within four rounds of hop, eliminating the need for an additional hop. We set $n_{hop}$ as 4 in Table 1 and 2. We also evaluate the answer performances as $n_{hop}$ varies and show the overall results in Appendix A.8.
Ablation on Traversal Model
In order to generalize HopRAG to scenarios with less computational overhead during retrieval, we supplement results from (1) HopRAG with traversal model Qwen2.5-1.5B-Instruct (2) HopRAG with non-LLM graph traversal that replaces the reasoning phase in Algorithm 1 with similarity matching. Table 5 shows that even without using the reasoning ability of LLM in the graph traversal, HopRAG can achieve 45.84% higher than BM25 and 25.43% higher than dense retriever (BGE), which proves the effectiveness of HopRAG in capturing textual similarity and logical relations for logic-aware retrieval. The introduction of reasoning ability from LLM (GPT-4o-mini) can achieve about 45.78% higher average score than the non-LLM version, and Qwen2.5-1.5B-Instruct as traversal model produces comparable results with less cost and higher efficiency. We analyze the retrieval efficiency in Appendix A.7.
@@ 5 Conclusion
Extensive experiments on multi-hop QA benchmarks, i.e. MuSiQue, 2WikiMultiHopQA, and HotpotQA, demonstrate that HopRAG outperforms conventional RAG systems and state-of-the-art baselines. Specifically, HopRAG achieved over 36.25% higher answer accuracy and 20.97% improved retrieval F1 score compared to conventional information retrieval approaches. It highlights the effectiveness of integrating logical reasoning into the retrieval module. Moreover, ablation studies provide insights into the sensitivity of hyperparameters and models, revealing trade-offs between retrieval performance and computational costs.
HopRAG paves the way toward reasoning-driven knowledge retrieval. Future work involves scaling HopRAG to broader domains beyond QA tasks; optimizing indexing and traversal strategies for more complex scenarios with lower computation costs.
Acknowledgments
This work is supported by the National Key R&D Program of China (2024YFA1014003), National Natural Science Foundation of China (92470121, 62402016), CAAI-Ant Group Research Fund, and High-performance Computing Platform of Peking University.
Limitations
@@ A.1 Symbols
$P$ passage corpus $Q_{i}^{+}$ set of out-coming questions for $p_{i}$
$p$ passage $q_{i,j}^{+}$ the j-th out-coming question for $p_{i}$
$q$ query $Q_{i}^{-}$ set of in-coming questions for $p_{i}$
$C$ retrieval context $q_{i,j}^{-}$ the j-th in-coming question for $p_{i}$
@@
$\mathcal{E}$ set of directed edges $\mathbf{v}_{i,j}^{+}$ embedding for $q_{i,j}^{+}$
@@
$k_{q}$ keywords for $q$ $\mathbf{v}_{i,j}^{-}$ embedding for $q_{i,j}^{-}$
$\mathbf{v}_{q}$ embedding for $q$ $R^{+}_{i}$ set of out-coming triplets
$C_{count}$ counter of vertices during traversal $r_{i,j}^{+}$ out-coming triplet for $q_{i,j}^{+}$
@@ A.3 Metrics
$F1_{\text{retrieval}}=\frac{2\cdot P\cdot R}{P+R}$
@@ A.4 Settings
To avoid semantic loss by chunking the documents at a fixed size, we chunk each document in a way corresponding to the supporting facts of each dataset. Specifically, we chunk each document in HotpotQA and 2WikiMultiHopQA by sentence since the smallest unit of these two datasets’ supporting facts is a sentence. To get embedding representation for each chunk, we use bge-base model. To extract keywords, we use the part-of-speech tagging function of Python package PaddleNLP to extract and filter entities. In our method, we use the Neo4j graph database to store vertices and build edges. When building edges we employ prompt engineering technique to instruct the LLM to generate an appropriate number of questions for each vertex to cover its information, with a minimum requirement of at least 2 in-coming questions and 4 out-coming questions. To prevent the graph structure from becoming overly complex and dense, we retain only $O(n\cdot\log(n))$ edges, where $n$ is the number of vertices. We use GPT-4o-mini for reasoning-augmented graph traversal, GPT-4o and GPT-3.5-turbo for inference with 2048 max tokens and 0.1 temperature in our main experiments.
For sparse and dense retrievers, we use the Neo4j database to conduct retrieval on the vertices. With this setting, we align the retrieval engine for unstructured baselines with HopRAG to fairly demonstrate the effectiveness of our graph structure index. For query decomposition, we use GPT-4o-mini to break down the query into multiple sub-queries, each of which should be a single-hop query. With $m$ sub-queries we conduct dense retrieval with BGE for each sub-query to get $top_{k}/m$ candidates independently and combine them all for final context. For reranking baseline, we use bge-reranker-base to rank $2*top_{k}$ candidates from dense retriever BGE and keep the $top_{k}$ ones as the final context. The structured baseline methods rely on specific open-source projects according to their papers.
@@ A.5 Prompts
(a) Demonstration of one edge between two vertices.
@@
(a) Demonstration of HopRAG’s graph.
(b) Demonstration of GraphRAG’s graph.
Figure 7: Visualizations of HopRAG and GraphRAG
@@ A.6 Case Study
We demonstrate the graph structure in Figure 7(a), one example edge with two vertices in Figure 6(a). Using the query "Donnie Smith who plays as a left back for New England Revolution belongs to what league featuring 22 teams?" as an example we conduct a qualitative analysis. For this multi-hop question (correct answer: Major League Soccer), the HotpotQA corpus contains three relevant sentences: (1) "Donald W. Donnie Smith (born December 7, 1990 in Detroit, Michigan) is an American soccer player who plays as a left back for New England Revolution in Major League Soccer."; (2) "Major League Soccer (MLS) is a men’s professional soccer league, sanctioned by U.S. Soccer, that represents the sport’s highest level in both the United States and Canada." and (3) "The league comprises 22 teams in the U.S. and 3 in Canada."
@@
Using this query and its passages for demonstration, we provide the visualizations of HopRAG’s graph index against GraphRAG to help readers grasp the differences and innovations. As shown in Figure 7 , the main differences between HopRAG and GraphRAG are listed as follows.
  • Vertices. GraphRAG uses LLM to summarize the information from text chunks and then creates additional vertices (e.g., event, organization and person) in the graph, while HopRAG directly stores the original chunks in the vertices and thus avoids LLM hallucination during summarization, information loss during entity extraction and overly dense graph structure from redundant vertices.
  • Edges. GraphRAG connects vertices with pre-defined relationships like "part of" or "related", while HopRAG flexibly stores in-coming questions on the edges along with their keywords and embeddings, which can not only guide reasoning-augmented graph traversal but also facilitate edge retrieval.
  • Index. GraphRAG creates and stores embeddings for the summarizations from LLM, while HopRAG creates sparse and dense indexes for both the vertices and the edges, which leads to more precise and efficient information retrieval.
In summary, the graph structure of HopRAG not only excavates logical relationships without creating additional vertices, but also paves the way for reasoning-driven knowledge retrieval.
@@
Table 8: We test the effect of hyperparameter $n_{hop}$ on HopRAG using GPT-3.5-turbo with top 20 passages. We vary $n_{hop}$ from 1 to 4 and report both the answer and retrieval metrics. For answer metrics, we report the answer EM and F1 score; For retrieval metrics, we report the F1 score and average number of calling LLM during traversal to measure the cost (the lower, the better). The best score is in bold and the second best is underlined.
A.7 Retrieval Efficiency
We supplement more comprehensive analysis of retrieval efficiency, along with optimization strategies for further speedup. The main latency in HopRAG’s retrieval comes from the LLM inference time during retrieval-augmented graph traversal. Since HopRAG with locally deployed Qwen2.5-1.5B-Instruct as the traversal model also showcases competitive performances, we focus on the retrieval efficiency in this scenario. Following the hyperparameters $n_{hop}=4$ and $top_{k}=20$ from our main experiments, each question requires calling LLM 38.53 times, where each LLM call involves selecting one edge from an average of 5.87 edges, with an input of around 500 tokens and an output of 20 tokens. According to (Qwen Development Team, 2025), the output token speed for Qwen2.5-1.5B-Instruct is about 40.86 token/s using BF16 and Transformer. Therefore, the additional latency for each question from LLM inference will be $38.53*20/40.86=18.86$ seconds. However, there are many optimization strategies to improve the retrieval efficiency. Using vLLM and GPTQ-Int4 techniques, the additional latency for each question can be reduced to $38.53*20/174.04=4.43$ seconds. Moreover, parallelism techniques like multithreading can further reduce the total execution time for all the queries.
A.8 Discussion Results on $n_{hop}$
Figure 8: Prompt for reasoning-augmented graph traversal.